OpenStudy (anonymous):

Solve the equation -4 + 6x = 17

OpenStudy (anonymous):

Are you just posting homework?

OpenStudy (anonymous):

no im studying

OpenStudy (anonymous):

To solve your equation, keep in mind that what you do on one side of (=), you have to do on the other.

OpenStudy (anonymous):

You want to "isolate" x and have x alone on one side.

OpenStudy (anonymous):

In this case, you add 4 to both sides first. That gives you: 6x = 17 + 4 Can you take it from here? Hint: Next step, divide both sides by 6 to isolate x.

OpenStudy (tansheet):

6x=17+4 6x=21 x=21/6 x=3.5 hence the value of x is 3.5